OpenStudy (anonymous):

Graph each pair of parametric equations. x=2t-1 y=t^2+5, -4 less than or equal to t less than or equal to 4

OpenStudy (anonymous):

x=2t-1 \[y=t ^{2}+5, -4 \le t \le 4\]

OpenStudy (dumbcow):

use substitution to get a function in terms of x \[t = \frac{x+1}{2}\] \[y = \frac{(x+1)^2}{4} +5\] this is a parabola with vertex at (-1,5) change the limits ---> -9< x < 7 graph: